# Forecasting Electricity Market Signals

### v4

@@ 4.3 LMP forecasting for Energy Market Participation
Table 3: Point evaluation of forecasting results for real-time price forecasting at Long Island. The numbers in the parentheses were the ranking of the algorithm. The columns under the label of LONGIL are the GPF performance of the 12-step foresting of the LMP at LONGIL. The columns under LONGIL & NYC are the performance of LMP forecasting at LONGIL using both LONGIL and NYC observations.
Methods NMSE NMAE MASE sMAPE
LONGIL LONGIL & NYC LONGIL LONGIL & NYC LONGIL LONGIL & NYC LONGIL LONGIL & NYC
SNARX [4] (8) $0.9852$ (7) $0.5029$ (8) $0.9733$ (8) $0.7318$ (8) $0.7053$ (8) $0.4548$ (7) $0.4319$ (7) $0.4162$
WIAE-GPF (1) $\mathbf{0.0585}$ (2) $0.0487$ (1) $\mathbf{0.2074}$ (1) $\mathbf{0.1186}$ (1) $\mathbf{0.1503}$ (1) $\mathbf{0.0737}$ (1) $\mathbf{0.0839}$ (1) $\mathbf{0.0316}$
TLAE [21] (4) $0.2956$ (1) $\mathbf{0.0232}$ (6) $0.4186$ (2) $0.1366$ (6) $0.3034$ (2) $0.0849$ (3) $0.2720$ (2) $0.3190$
DeepVAR [11] (6) $0.3919$ (6) $0.4060$ (5) $0.4088$ (7) $0.4097$ (5) $0.2963$ (7) $0.2546$ (5) $0.3629$ (3) $0.3703$
LQRA [16] (7) $0.8287$ (8) $0.7110$ (7) $0.8472$ (4) $0.2693$ (7) $0.6139$ (4) $0.1623$ (8) $0.4952$ (8) $0.4879$
BWGVT [19] (3) $0.2670$ (5) $0.2528$ (4) $0.3158$ (6) $0.3280$ (4) $0.2289$ (6) $0.2038$ (4) $0.2817$ (6) $0.3900$
Pyraformer [34] (5) $0.3128$ (3) $0.1382$ (2) $0.3074$ (3) $0.2556$ (2) $0.2228$ (3) $0.1588$ (2) $0.1059$ (4) $0.3765$
Informer [33] (2) $0.1912$ (4) $0.1829$ (3) $0.3147$ (5) $0.2830$ (3) $0.2281$ (5) $0.1759$ (6) $0.3761$ (5) $0.3827$
Table 4: Evaluation of probabilistic forecasting results of real-time prices at Long Island. The numbers inside the square parentheses are the corresponding NCW.
Methods CRPS CPE (90%) [NCW] CPE (50%) [NCW] CPE (10%) [NCW]
LONGIL LONGIL & NYC LONGIL LONGIL & NYC LONGLIL LONGIL & NYC LONGLIL LONGIL & NYC
SNARX [4] (3) $18.2864$ (3) $10.5286$ (5) $-0.2263$ $[0.6620]$ (6) $-0.3779$ $[0.7990]$ (5) $-0.2451$ $[0.5574]$ (2) $-0.0746$ $[0.9161]$ (3) $-0.0641$ $[0.8145]$ (6) $0.1870$ $[0.8082]$
WIAE-GPF (1) $\mathbf{4.6029}$ (1) $\mathbf{1.1519}$ (1) ${}\mathbf{0.0245}$ $[0.9936]$ (1) ${}\mathbf{0.0157}$ $[0.9958]$ (1) $\mathbf{0.0216}$ $[0.9681]$ (1) $\mathbf{0.0184}$ $[0.8742]$ (1) $\mathbf{0.0242}$ $[0.9836]$ (1) $\mathbf{-0.0190}$ $[0.8207]$
TLAE [21] (2) $6.1874$ (2) 2.84492.8449{2.8449}2.8449 (2) 0.03250.0325~{}0.03250.0325 $[0.8179]$ (2) 0.05920.0592~{}0.05920.0592 $[0.8176]$ (2) $0.0423$ $[0.6357]$ (3) $0.1286$ $[0.7637]$ (5) $-0.0848$ $[0.6704]$ (4) $-0.0834$ $[0.6516]$
DeepVAR [11] (4) $23.1460$ (5) $22.9343$ (4) $-0.0575$ $[0.9738]$ (3) $-0.0761$ $[0.8467]$ (4) $-0.1637$ $[0.5709]$ (5) $-0.1488$ $[0.6614]$ (2) $-0.0616$ $[0.5889]$ (3) $-0.0431$ $[0.4260]$
LQRA [16] (6) $48.9460$ (4) $17.0583$ (6) $-0.2601$ $[0.7130]$ (5) $-0.0958$ $[0.8434]$ (3) $-0.0465$ $[0.6534]$ (4) $0.1364$ $[0.9082]$ (4) $0.0648$ $[0.8494]$ (5) $0.1529$ $[0.8827]$
BWGVT [19] (5) $24.2595$ (6) $24.3065$ (3) 0.09990.0999~{}0.09990.0999 $[7.3500]$ (4) 0.08290.0829~{}0.08290.0829 $[3.4864]$ (6) $0.4815$ $[2.7822]$ (6) $0.4322$ $[3.0100]$ (6) $0.1034$ $[7.2325]$ (2) $0.0311$ $[1.1181]$
@@
Figure 5: Trajectories of the real-time price at LONGIL, and its prediction generated by selective methods.
The real-time market of NYISO closes sixty minutes ahead of actual delivery, which means that the forecasting horizon needs to be longer than 60 minutes. Two experiments were conducted to produce probabilistic forecasts of 60-minute ahead LMPs at the Long Island (LONGIL) using (a) the day-ahead prices and the current and past real-time LMPs at LONGIL, along with the system load up to the time of submitting the bid; (b) the neighboring NYC real-time LMP in addition to the data in (a).
Fig. 4 shows the real-time LMP trajectories at both LONGIL and NYC along with the demand and the day-ahead LMP at LONGIL. The real-time LMPs at LONGIL and NYC showed apparent spatial dependencies, while the dependency between day-ahead and real-time LMPs at LONGIL. The dependencies between load and real-time LMP were less obvious. The real-time LMPs and load were collected every 5 minutes and day-ahead LMPs every hour. We used the first 25 days for training and validation, and last 6 days of July for evaluation.
Test results are shown in Table 3 and 4 with boldfaced numbers being the best performance. We observed that WIAE-GPF performed the best for all cases except for the forecasting of LONGIL LMP using both LONGIL and NYC data, for which WIAE-GPF is close second under NMSE. The strong performance of WIAE-GPF comes from that it aims to match conditional distribution, with a validity guarantee that ensures the Monte-Carlo samples generated have the same conditional distribution as that of the actual time series variable. Overall, the second-best is TLAE; the VAE-trained RNN autoencoder achieved noticeable gain by combining data from two locations, especially for the joint forecasting at LONGIL & NYC. However, under the probabilistic forecasting measure of CRPS, TLAE was 50% to 100% higher than that of WIAE-GPF. Because sampling the correlated Gaussian latent process is nontrivial, TLAE used a re-parameterization heuristic that could cause an accumulation of biases.
The three LLM-based forecasters (BWGVT, Pyraformer, and Informer) performed roughly the same, worse than WIAE-GPF and TLAE but better than SNARX and DeepVAR. Note that Pyraformer and Informer are both point forecasters trained to minimize the mean squared forecasting error. They did not perform well under NMAE, MASE, and sMAPE. BWGVT is adapted to be a probabilistic forecaster, although it did not perform well under the CRPS score. Its quantile prediction was outlier-sensitive, especially when compared with the GPF methods that depend on the stochastic latent process. We also observed that BWGVT predicted larger intervals as shown by its high NCW for all cases. Consequently, their CPE’s were always positive, indicating that the prediction intervals it generated covered more than the nominal percentages. Hence, both their point estimation results and CRPS were worse than TLAE and WIAE-GPF.
The other quantile regression based method, LQRA, performed worst among nonparametric techniques. Originally proposed for hourly price forecasting, LQRA’s assumed an simple regression model to produce point forecasts, with nonlinearity only introduced by the one-day ahead minimum and maximum day-ahead price. Although this model might be sufficient for day-ahead price forecasting, its model mismatch when applied to real-time price forecasting intensifies the inaccuracy.
For the two parametric techniques DeepVAR and SNARX, being significantly simpler than LLM forecasters, performed slightly worse. Both techniques did not perform well under point estimation metrics. Shown by the CPE’s in Table 4, their prediction interval estimation is also worse than other probabilistic forecasting methods. Their NCW showed that they predicted narrow intervals that failed to cover the majority of the ground truths. The inaccurate semiparametric-AR model assumption appears to be culpable. Their performance was influenced by model mismatch, where ARMA models typically expect smoother trajectories. For the real-time LMP that exhibits high volatility, these methods were slow catch the rapid changes, leading to predicting shifted peaks that resulted in huge forecasting errors. This phenomena were corroborated by their high sMAPE, indicating higher percentage error at each time stamp.
It is interesting to observe that using LMPs in neighboring locations improved the forecasting performance of all methods except DeepVAR and BWGVT, as shown by the LONGIL and LONGIL & NYC columns under the DeepVAR and BWGVT rows in Table 3. This implies that these two methods didn’t account for spatial correlations optimally. We conclude this to the difficulty of training of transformer and LSTM models.
@@
To gain insights into the performance of WIAE-GPF and other benchmark techniques, we plotted the ground truth trajectories (black) and trajectory forecasts generated by WIAE-GPF (red) and a competing algorithm (blue) in Fig. 5. Note that the spikes were not predicted by any methods. This was not surprising given the nature of how these spikes were produced. Aside from these spikes, these figures show clearly that WIAE-GPF (red) tracked the ground truth (black) the closest, which was supported by the fact that WIAE-GPF has the smallest NMAE. We also observed that WIAE-GPF had the smallest variation, which is supported by the fact that WIAE-GPF had the smallest NMSE. Furthermore, WIAE-GPF was the least affected by the price spikes. This was because, as a GPF method, the Monte-Carlo samples used to produce the MMSE point estimate were less likely to include extreme samples. However, for SNARX, as shown in Fig. 5(c), the spikes caused significant deviations from the forecasted trajectory.
@@ 4.4 Interregional LMP spread for Interchange Markets
Methods NMSE NMAE MASE sMAPE PER
SNARX [4] (7) $2.4531$ (7) $1.3415$ (7) $1.1847$ (7) $0.4958$ (7) $0.6781$
WIAE-GPF (1) $\mathbf{0.0098}$ (1) $\mathbf{0.2738}$ (1) $\mathbf{0.2418}$ (1) $\mathbf{0.4493}$ (1) $\mathbf{0.0606}$
TLAE [21] (5) $0.9592$ (5) $0.9785$ (5) $0.8641$ (4) $0.4785$ (4) $0.3692$
DeepVAR [11] (6) $1.8986$ (3) $0.7224$ (3) $0.6380$ (5) $0.4806$ (3) $0.3505$
BWGVT [19] (3) $0.9053$ (4) $0.8525$ (4) $0.7529$ (3) $0.4674$ (2) $0.2313$
Pyraformer [34] (4) $0.9478$ (6) $1.2674$ (6) $1.1193$ (6) $0.4909$ (6) $0.6738$
Informer [33] (2) $0.8045$ (2) $0.4185$ (2) $0.4252$ (2) $0.4580$ (5) $0.5487$
Table 6: Probabilistic forecasting results of spread forecasting between NYISO and PJM.
Methods CRPS CPE (90%) [NCW] CPE (50%) [NCW] CPE (10%) [NCW]
SNARX [4] (5) $120.0403$ (5) $-0.5443$ $[5.0419]$ (5) $-0.2958$ $[1.4555]$ (2) $-0.0195$ $[0.6015]$
WIAE-GPF (1) $\mathbf{4.0329}$ (1) $\mathbf{0.0215}$ $[0.4427]$ (2) ${-0.0274}$ $[0.5255]$ (3) $0.0212$ $[0.1692]$
TLAE [21] (2) $15.5195$ (3) $-0.0443$ $[0.7745]$ (1) $\mathbf{0.0052}$ $[0.8784]$ (5) $-0.0933$ $[0.3133]$
DeepVAR [11] (4) $32.8296$ (2) $-0.0355$ $[2.0279]$ (3) ${-0.1480}$ $[1.4739]$ (1) $\mathbf{-0.0021}$ $[0.4198]$
BWGVT [19] (3) $31.5660$ (4) $0.0989$ $[5.1788]$ (4) $0.1835$ $[6.0939]$ (4) $0.0656$ $[4.0473]$
@@
Seen from Table 5 and 6, WIAE-GPF performed better than all other techniques in all metrics. TLAE performed the second-best in CRPS ($15.5195$) but slightly worse than BWGVT when evaluated under point estimation metrics. Its sequential sampling of the latent Gaussian process added to its numerical instability. BWGVT was the overall second-best performing probabilistic technique. Its transformer architecture with enhanced capability of capturing long-term temporal dependency didn’t offer much gain for the training difficulty imposed by the increasing number of deep-learning parameters, see Sec. 4.6. BWGVT also exhibited the tendency to predict a wide interval covering more than the nominal percentage. Point estimation techniques, namely Pyraformer and Informer, were not competitive when evaluated under point estimation metrics other than NMSE. Among probabilistic methods, DeepVAR performed similarly to the LLM methods, and SNARX had the most difficulties. These (semi) parametric methods suffered from model mismatch, and were sensitive to sudden changes, Thus, shifted peaks and valleys were often witnessed in their predictions.
Same observation can also be made through Fig. 6. WIAE-GPF has the most stable prediction of interregional LMP spreads, which is corroborated by its smallest NMSE and NMAE. Pyraformer also follows the trend of LMP spreads accurately but with higher variance. The two AR-based parametric models, SNARX and DeepVAR exhibited the tendency to predict shifted spikes and failures to catch the rapid and dramatic change of LMP spread.

### v5

@@ 4.2 LMP forecasting for Energy Market Participation
Table 4: Performance evaluation of forecasting results for real-time price forecasting at Long Island. The numbers in the parentheses were the ranking of the algorithm. The columns under the label of LONGIL are the GPF performance of the 12-step foresting of the LMP at LONGIL.
Season Methods NMSE NMAE MASE sMAPE CRPS CPE (90%) [NCW] CPE (50%) [NCW] CPE (10%) [NCW]
Summer (2023.07/01-2023.07.31) WIAE-GPF (1) $\mathbf{0.1084}$ (2) $0.2341$ (2) $0.5812$ (2) $0.2644$ (1) $\mathbf{13.2607}$ (1) $\mathbf{0.0009}[0.0718]$ (1) $\mathbf{-0.0264}[0.0261]$ (1) $\mathbf{0.0059}[0.0055]$
TLAE [22] (6) $0.4690$ (6) $0.4616$ (6) $0.9768$ (6) $0.4436$ (4) $23.3288$ (4) $-0.1034[0.0723]$ (3) $0.1492[0.0268]$ (3) $-0.0646[0.0027]$
DeepVAR [14] (3) $0.2009$ (1) $\mathbf{0.1849}$ (1) $\mathbf{0.3911}$ (1) $\mathbf{0.2496}$ (2) $14.0716$ (2) $0.0275[0.0777]$ (2) $-0.0459[0.226]$ (2) $-0.0243[0.0034]$
BWGVT [28] (4) $0.2206$ (4) $0.2804$ (4) $0.5934$ (3) $0.2711$ (3) $15.0476$ (3) $0.0968[0.0862]$ (4) $0.1646[0.0419]$ (4) $0.1279[0.0081]$
Pyraformer [30] (2) $0.1090$ (3) $0.2717$ (3) $0.5749$ (4) $0.2934$ N/A N/A N/A N/A
Informer [31] (5) $0.4505$ (5) $0.3935$ (5) $0.8326$ (5) $0.3329$ N/A N/A N/A N/A
Fall (2023.10.01-2023.10.31) WIAE-GPF (1) $\mathbf{0.1020}$ (2) $0.2806$ (2) $0.4131$ (2) $0.2522$ (1) $\mathbf{10.9362}$ (1) $\mathbf{0.0002}[0.0711]$ (1) $\mathbf{-0.0390}[0.0292]$ (1) $\mathbf{-0.0157}[0.0054]$
TLAE [22] (6) $0.6181$ (6) $0.4693$ (6) $0.6908$ (6) $0.4349$ (4) $19.1394$ (3) $-0.0487[0.0335]$ (3) $-0.0395[0.0135]$ (2) $0.0230[0.0059]$
DeepVAR [14] (5) $0.5684$ (1) $\mathbf{0.2328}$ (1) $\mathbf{0.3427}$ (3) $0.2662$ (2) $12.6109$ (2) $0.0205[0.0414]$ (2) $-0.0328[0.106]$ (3) $-0.0246[0.0018]$
BWGVT [28] (4) $0.2992$ (5) $0.3696$ (5) $0.5441$ (5) $0.3098$ (3) $14.1423$ (4) $0.0905[0.0837]$ (4) $0.1763[0.0393]$ (4) $0.0367[0.0074]$
Pyraformer [30] (2) $0.1080$ (3) $0.2949$ (3) $0.4341$ (1) $\mathbf{0.2116}$ N/A N/A N/A N/A
Informer [31] (3) $0.2615$ (4) $0.2985$ (4) $0.4394$ (4) $0.2823$ N/A N/A N/A N/A
Winter (2024.01.01-2024.01.31) WIAE-GPF (2) 0.09060.0906{0.0906}0.0906 (2) $0.2263$ (2) $0.3369$ (1) $\mathbf{0.2307}$ (1) $\mathbf{12.1949}$ (1) $\mathbf{0.0014}[0.1203]$ (2) ${0.0358}[0.0493]$ (2) ${0.0196}[0.0092]$
TLAE [22] (6) $0.2244$ (6) $0.3306$ (6) $0.4922$ (6) $0.4355$ (4) $20.7468$ (3) $-0.0337[0.0371]$ (3) ${-0.0369}[0.0159]$ (4) $-0.0502[0.0027]$
DeepVAR [14] (3) $0.1507$ (1) $\mathbf{0.1614}$ (1) $\mathbf{0.3273}$ (2) $0.2499$ (2) $13.1403$ (2) $0.0211[0.1525]$ (1) $\mathbf{-0.0160}[0.220]$ (3) $-0.0287[0.0049]$
BWGVT [28] (5) $0.1649$ (4) $0.2725$ (4) $0.4057$ (3) $0.2515$ (3) $15.2919$ (4) $0.081[0.1808]$ (4) $0.1804[0.0545]$ (1) $\mathbf{0.0158}[0.0058]$
Pyraformer [30] (1) $\mathbf{0.0721}$ (5) $0.2786$ (5) $0.4144$ (5) $0.2747$ N/A N/A N/A N/A
Informer [31] (4) $0.1553$ (3) $0.2528$ (3) $0.3763$ (4) $0.2617$ N/A N/A N/A N/A
Spring (2024.04.01-2024.04.30) WIAE-GPF (1) $\mathbf{0.0797}$ (2) $0.2137$ (2) $0.2942$ (2) 0.12890.1289{0.1289}0.1289 (1) $\mathbf{5.3111}$ (1) $\mathbf{0.0068}[0.1485]$ (1) $\mathbf{0.0208}[0.0624]$ (1) $\mathbf{0.0148}[0.0113]$
TLAE [22] (6) $0.1903$ (6) $0.2729$ (6) $0.3756$ (6) $0.4016$ (4) $15.3330$ (3) $-0.0510[0.1270]$ (2) ${0.0273}[0.0635]$ (2) $0.0217[0.0143]$
DeepVAR [14] (5) $0.1105$ (4) 0.23450.2345{0.2345}0.2345 (4) 0.32280.3228{0.3228}0.3228 (1) $\mathbf{0.1233}$ (2) $5.9420$ (2) $0.0154[0.1665]$ (3) $-0.0423[0.572]$ (3) $-0.0415[0.0069]$
BWGVT [28] (3) $0.0877$ (3) $0.2188$ (3) $0.3011$ (5) $0.2166$ (3) $14.6103$ (4) $0.0805[0.1551]$ (4) $0.2436[0.0871]$ (4) ${0.1177}[0.0196]$
Pyraformer [30] (4) $0.1022$ (5) $0.2890$ (5) $0.3978$ (3) $0.1788$ N/A N/A N/A N/A
Informer [31] (2) $0.0845$ (1) $\mathbf{0.1867}$ (1) $\mathbf{0.2570}$ (4) $0.1880$ N/A N/A N/A N/A
Summer real-time LMP prediction
Test results for July 2023 are shown in row 2-7 of Table 4 with the best performances highlighted in bold. We observed that WIAE-GPF performed the best for the probabilistic evluation metrics and NMSE, and ranked in close second for NMAE, MASE and sMAPE, whereas DeepVAR performed the best for those three metrics associated with conditional median. Pyraformer, trained to minimize MSE, achieved the second-best NMSE but didn’t achieve notable performance for the other metrics. We observed that WIAE-GPF outperformed other methods in probabilistic evaluation metrics and NMSE, while ranking a close second in NMAE, MASE, and sMAPE. In contrast, DeepVAR excelled in these three metrics, which are associated with the conditional median. Pyraformer, which was trained to minimize MSE, achieved the second-best NMSE but did not show notable performance in the other metrics. We believe this may be due to DeepVAR, as a Gaussian AR process with fewer parameters and a simpler training objective, could perform better at estimating the conditional median. On the other hand, WIAE-GPF, being nonparametric and trained with two Wasserstein discriminators, is more sensitive to hyperparameters and initialization. However, its nonparametric nature allows it to avoid the model misspecification issues that can affect parametric methods like DeepVAR when predicting the overall conditional distribution, as evidenced by its superior performance in CRPS and CPE.
@@
Figure 5: Trajectories of the real-time price at LONGIL in July (top row) and October (bottom row) 2023, and its prediction generated by selective methods.
(a) WIAE-GPF
(b) Pyraformer
(c) TLAE
(d) DeepVAR
(e) WIAE-GPF
(f) Informer
(g) TLAE
(h) DeepVAR
Figure 6: Trajectories of the real-time price at LONGIL in January (top row) and April 2024 (bottom row), and its prediction generated by selective methods.
@@
To gain insights into the performance of WIAE-GPF and other benchmark techniques, we plotted the ground truth trajectories (black) and trajectory forecasts generated by WIAE-GPF (red) and a competing algorithm (blue) in Fig. 5 (top row). Note that the spikes were not predicted by any methods. This was not surprising given the nature of how these spikes were produced. Aside from these spikes, these figures show clearly that WIAE-GPF (red) tracked the ground truth (black) the closest, which was supported by the fact that WIAE-GPF has the smallest NMAE. We also observed that WIAE-GPF had the smallest variation, which is supported by the fact that WIAE-GPF had the smallest NMSE. Furthermore, WIAE-GPF was the least affected by the price spikes. This was because, as a GPF method, the Monte-Carlo samples used to produce the MMSE point estimate were less likely to include extreme samples.
Fall Real-time LMP Prediction
Test results for October are shown in row 8-13 of Table 4 displayed similarity to summer’s results. The volatility of fall LMPs is slightly lower than the volatility of summer. WIAE-GPF performed the best for all except for NMAE, MASE and sMAPE, where it ranked the second. DeepVAR performed the best for NMAE and MASE, where conditional median is used as the point forecastor. Pyraformer achieved the best sMAPE and the second-best NMSE.
The performance can also be understood by examining the trajectories shown in Fig. 5 (bottom row). Similarly, WIAE-GPF exhibited the smallest variation and remained closest to the ground truth (excluding the spikes). In contrast, Fig. 5(h) reveals that DeepVAR tended to predict shifted peaks. This behavior is characteristic of the AR process, which is heavily influenced by past observations.
Winter Real-time LMP Prediction
January 2024 has the most volatile LMP among the four months tested. Seen from Table. 4, WIAE-GPF obtained the best sMAPE, CRPS, and CPE(90%), and came close in the second place for the rest. Pyraformer, with its MSE-minimizing training objective, achieved the best NMSE at around $7\%$. DeepVAR had the best NMAE, MASE and CPE(50%), proving its capability of accurately estimating conditional distribution around the median.
Fig. 6 (top row) showed the trajectory of winter LMP predictions. DeepVAR exhibited larger variability when predicting LMP of January 2024, which explains its higher level of NMSE. Pyraformer and WIAE-GPF achieved similar level of variance when predicting winter LMP. TLAE had the worst ground truth-tracking performance among the four trajectories.
Spring Real-time LMP Prediction
April 2024 has the least volatile LMPs among the four months. For this rolling-window experiment, WIAE-GPF achieved the best metrics except for NMAE, MASE and sMAPE. Informer performed better than Pyraformer on the spring dataset, with the best NMAE and MASE. DeepVAR obtained the best sMAPE.
Fig. 6 (bottom row) showed the trajectory of spring LMP predictions. DeepVAR and TLAE exhibited larger deviation for the spring , which explains its higher level of NMSE. Informer exhibited the best ground-truth tracking capability, with WIAE-GPF came in close second.
@@ 4.3 Interregional LMP spread for Interchange Markets
Methods NMSE NMAE MASE sMAPE PER CRPS CPE (90%) [NCW] CPE (50%) [NCW] CPE (10%) [NCW]
WIAE-GPF (1) $\mathbf{0.0098}$ (1) $\mathbf{0.2738}$ (1) $\mathbf{0.2418}$ (1) $\mathbf{0.4493}$ (1) $\mathbf{0.0606}$ (1) $\mathbf{4.0329}$ (1) $\mathbf{0.0215}$ $[0.4427]$ (2) ${-0.0274}$ $[0.5255]$ (3) $0.0212$ $[0.1692]$
TLAE [22] (5) $0.9592$ (5) $0.9785$ (5) $0.8641$ (4) $0.4785$ (4) $0.3692$ (2) $15.5195$ (3) $-0.0443$ $[0.7745]$ (1) $\mathbf{0.0052}$ $[0.8784]$ (5) $-0.0933$ $[0.3133]$
DeepVAR [14] (6) $1.8986$ (3) $0.7224$ (3) $0.6380$ (5) $0.4806$ (3) $0.3505$ (4) $32.8296$ (2) $-0.0355$ $[2.0279]$ (3) ${-0.1480}$ $[1.4739]$ (1) $\mathbf{-0.0021}$ $[0.4198]$
BWGVT [28] (3) $0.9053$ (4) $0.8525$ (4) $0.7529$ (3) $0.4674$ (2) $0.2313$ (3) $31.5660$ (4) $0.0989$ $[5.1788]$ (4) $0.1835$ $[6.0939]$ (4) $0.0656$ $[4.0473]$
Pyraformer [30] (4) $0.9478$ (6) $1.2674$ (6) $1.1193$ (6) $0.4909$ (6) $0.6738$ N/A N/A N/A N/A
Informer [31] (2) $0.8045$ (2) $0.4185$ (2) $0.4252$ (2) $0.4580$ (5) $0.5487$ N/A N/A N/A N/A
Methods NMSE NMAE MASE sMAPE CRPS CPE (90%) [NCW] CPE (50%) [NCW] CPE (10%) [NCW]
WIAE-GPF (1) $\mathbf{0.5957}$ (1) $\mathbf{0.7555}$ (1) $\mathbf{0.4698}$ (1) $\mathbf{0.1059}$ (1) $\mathbf{0.0081}$ (1) $\mathbf{-0.0016}$ $[0.9199]$ (1) $\mathbf{0.0321}$ $[0.9336]$ (1) $\mathbf{-0.0132}$ $[0.8885]$
TLAE [22] (5) $1.1727$ (5) $1.0605$ (5) $0.6595$ (3) $0.2782$ (4) $1.5541$ (4) $-0.7857$ $[0.0004]$ (4) $-0.4489$ $[0.0005]$ (4) $-0.0957$ $[0.0027]$
DeepVAR [14] (7) $1.4431$ (7) $1.1750$ (7) $0.7307$ (5) $0.3952$ (3) $1.2947$ (3) $-0.3526$ $[0.5665]$ (3) $-0.2560$ $[0.5296]$ (3) $-0.0521$ $[0.5434]$
BWGVT [28] (3) $0.9562$ (2) $0.9793$ (2) $0.6090$ (4) $0.3168$ (2) $1.2488$ (2) $0.0065$ $[1.8309]$ (2) $0.0754$ $[2.0385]$ (5) $0.0996$ $[2.4261]$
Pyraformer [30] (4) $0.9783$ (4) $0.9948$ (4) $0.6186$ (7) $0.4986$ N/A N/A N/A N/A
Informer [31] (2) $0.6006$ (3) $0.9819$ (3) $0.6106$ (2) $0.2247$ N/A N/A N/A N/A
@@
Seen from Table 6, WIAE-GPF performed better than all other techniques in all metrics. TLAE performed the second-best in CRPS ($15.5195$) but slightly worse than BWGVT when evaluated under point estimation metrics. Its sequential sampling of the latent Gaussian process added to its numerical instability. BWGVT was the overall second-best performing probabilistic technique. Its transformer architecture with enhanced capability of capturing long-term temporal dependency didn’t offer much gain for the training difficulty imposed by the increasing number of deep-learning parameters, see Sec. 4.5. BWGVT also exhibited the tendency to predict a wide interval covering more than the nominal percentage. Point estimation techniques, namely Pyraformer and Informer, were not competitive when evaluated under point estimation metrics other than NMSE. Among probabilistic methods, DeepVAR performed similarly to the LLM methods. The (semi) parametric methods suffered from model mismatch, and were sensitive to sudden changes, Thus, shifted peaks and valleys were often witnessed in their predictions.
Same observation can also be made through Fig. 7. WIAE-GPF has the most stable prediction of interregional LMP spreads, which is corroborated by its smallest NMSE and NMAE. Pyraformer also follows the trend of LMP spreads accurately but with higher variance. The AR-based parametric models, DeepVAR exhibited the tendency to predict shifted spikes and failures to catch the rapid and dramatic change of LMP spread.